# SynSUM – Synthetic Benchmark with Structured and Unstructured Medical Records

### v2

@@ 2.2 Generating unstructured text with a large language model
Once the tabular patient record has been generated, we prompt a large language model (LLM, in our case GPT-4o) to write a clinical note based on the tabular variables associated with this fictional encounter. For this, we use a prompt which is generally structured like the example shown in Figure 1 (except for some special cases, which are addressed in Appendix C). Additional example prompts can be found in Appendix D. Only the background variables and symptoms may be directly mentioned in the prompt, while the diagnoses pneumonia and common cold would not yet be known to a clinician who is taking notes during a consultation and are therefore left out. They can still influence the content of the note through the descriptions of the symptoms that are included in the prompt. The treatment and outcome are left out of the prompt as well, just like the non-clinical variables, as all of these are typically either unknown or irrelevant at the time of writing the note.We ask the LLM to generate both a “history” portion and a “physical examination” portion, to encourage variation and complexity of the generated notes.
In general, the prompt is constructed as follows. For more details, we refer to Appendix C. First, we list the symptoms which are experienced by the patient. We do not exhaustively list the full set of symptoms, but rather sample the probability that a symptom is mentioned according to a distribution defined by the expert. This renders the notes more realistic, since real notes do not exhaustively mention the presence or absence of all possible symptoms, but rather follow the narrative of the patient and the subsequent probing of the clinician. While we encourage the LLM to invent additional context around the patient’s symptoms, we want this to at least partially relate to the underlying cause of the symptom. For this reason, we include a descriptor next to each symptom, which is randomly sampled from a set of expert-defined phrases that describe the symptom in the case where it results from a particular cause. This way, the diagnoses can influence the content of the note, even when they are not mentioned explicitly in the prompt. The full list of descriptors, as well as further information on how we sample them, is included in Appendix C. The next part of the prompt lists the underlying health conditions, which are assumed to be known up-front, and could therefore influence the content of the note. Finally, the prompt lists some additional instructions, which we motivate in Appendix C. There, we also outline an alternative prompting strategy, which is used to generate clinical notes for the patient records where none of the respiratory symptoms are present, which is the case for around one third of our dataset. In these cases, we encourage the LLM to imagine an alternative reason for the patient visit, unrelated to the respiratory domain.
@@
Real clinical notes can be challenging, often containing abbreviations, shortcuts and denser sentence structure. To make the notes more challenging, we ask the LLM to create a compact version of each note through an additional prompt, as can be seen in Figure 1. Our dataset contains both the original note and the compact version of the note.
3 Expert evaluation
Normal Compact
mean 4.69 4.53 4.15 4.92 4.88 4.02
std 0.12 0.21 0.30 0.07 0.10 0.31
@@ 3 Expert evaluation
    Realism: The context and details invented by the LLM and added to the “history” portion of the note should be realistic given the symptoms experienced by the patient, as detailed in the prompt. Furthermore, the elements that are checked in the “physical examination” must be realistic in light of the patient information described in the “history”. We evaluate both aspects separately, focusing on content (what information is included in the note), rather than format (how the information is written down).
@@
Our artificial notes were rated as highly consistent with the prompt, and therefore with the information present in the tabular portion of the dataset. As shown in Appendix E.2, a large majority of the inconsistencies arise from a violation of the additional instructions (usually by inventing additional symptoms), while the key information included in the prompt was still conveyed correctly in the note.
@@
4 Symptom predictor baselines
In order to set a baseline for future information extraction tasks, we run various prediction models on both the tabular and textual parts of the dataset. These models are trained to predict each of the five symptoms: dyspnea, cough, pain, fever and nasal.
Two of our baselines only get to see the tabular portion of the dataset at the input: Bayesian network (BN-tab) and XGBoost (XGBoost-tab). We use these models to predict each symptom in three settings, differing from one another in the set of tabular features that are taken as an input, which we call the evidence:
@@ 4 Symptom predictor baselines
Apart from the tabular-only baselines, we also train some baselines that get to see the text. Our neural-text classifier takes only the text as an input (in the form of a pretrained clinical sentence embedding) and outputs the probability that a symptom is mentioned in the text. We extend this text-only baseline by concatenating a numerical representation of the tabular features to the text embedding at the input, forming the neural-text-tab baseline. Again, we do this for each of the three evidence settings outlined above. Note that this is the only model that combines both the background knowledge available in the tabular features with the unstructured text, and it does so in a naive way. Future work will focus on improving the performance of this model by exploiting the relations between any of the tabular concepts, as envisioned in Figure 2.
@@
- normal + all 0.9526 0.9481 0.8096 0.9598 0.8804
- normal + no-sympt 0.9592 0.9530 0.8078 0.9550 0.9014
- normal + realistic 0.9544 0.9543 0.8303 0.9575 0.9101
@@
- compact + all 0.9535 0.9384 0.7675 0.9566 0.8987
- compact + no-sympt 0.9363 0.9240 0.7984 0.9638 0.8991
- compact + realistic 0.9442 0.9422 0.7880 0.9606 0.9051
Table 2: F1-score obtained over the test set for each of our baseline models. The results for the text classifiers trained over the normal vs. the compact version of the notes are grouped together for readability. We report results for the mean embedding type, while results for the other embedding types can be found in Table 8. We underline the best result obtained by the tabular-only models, while the best overall result per symptom is in shown in bold.
4.1 Models
BN-tab
We provide the causal structure in Figure 1 to the BN, and learn all parameters in the conditional probability tables (CPTs), Noisy-OR distributions, logistic regression model and Poisson regression model from the training data. In each case, we use maximum likelihood estimation to estimate the parameters. Where we don’t directly learn a CPT (for the variables dyspnea, cough, pain, nasal, antibiotics and #days), we evaluate the learned distribution for each combination of child and parent values to obtain a CPT. For more details, we refer to Appendix F.1. We then use variable elimination over the full joint distribution to evaluate the capability of the learned BN to predict each of the symptoms, taking different variables as evidence according to the three settings described earlier (all, no-sympt and realistic).
XGBoost-tab
We train an XGBoost classifier for each symptom in combination with each of the three evidence settings, meaning each classifier sees a different set of tabular features at the input. We optimize the hyperparameters separately for each combination (15 in total) using 5-fold cross-validation. For more details, we refer to Appendix F.2.
Neural-text
We train a neural classifier that takes only the text as an input and is trained to predict the probability a symptom is mentioned. We train separate classifiers for each symptom. We first split the text into sentences, and transform these into an embedding using the pretrained clinical representation model BioLORD-2023 (Remy, Demuynck, and Demeester 2024). We explore 4 settings for turning these sentence embeddings into a single note embedding:
The note embedding is then fed into a multi-layer perceptron with one hidden layer, followed by a Sigmoid activation for the symptoms dyspnea, cough, pain and nasal, and a Softmax activation with 3 outputs heads for the symptom $fever$. We optimized the parameters of each model (i.e. each combination of symptom and embedding type) using the binary or multiclass cross-entropy objective over the symptom labels. For more details, we refer to Appendix F.3.
Neural-text-tab
We extend the neural-text baseline by concatenating the mean text embeddings with the tabular variables at the input of each neural classifier. All categorical tabular variables were first transformed to a one-hot encoding, while the variable #days was preprocessed using standard scaling. We used the same architecture as the neural-text baseline (only changing the dimension of the input layer), and again trained separate classifiers for each symptom combined with each evidence setting (all, no-sympt and realistic). For more details, we refer to Appendix F.4.
4.2 Results
We use a random $8{,}000$/$2{,}000$ split to obtain a train and test set. We use cross-validation on the train set to tune any hyperparameters, and report the final F1-score over the test set after training. For the binary symptoms, we use a $0.5$ decision threshold. For fever, which has three possible categories, the class with the highest predicted probability is chosen. In that case, we report the macro F1-score.
@@ 4.2 Results
Table 2 compares the results obtained for all baselines. The tabular-only baselines (BN-tab and XGBoost-tab) perform consistently worse than the baselines that include text (neural-text and neural-text-tab). The evidence setting where all other features are included as evidence usually performs best for the tabular-only baselines.
@@
The neural-text-tab baseline does not perform better than the neural-text baseline for the normal notes. While there is little room for improvement in the dyspnea, cough and nasal classifiers, the symptoms pain and fever are harder to predict. We also note a consistent gap in performance between the normal and compact notes, which can be attributed to the higher complexity of the latter. In that case, the neural-text-tab classifier manages to marginally improve over the neural-text classifier by including the tabular features.
@@
Table 8 in Appendix F.3 further breaks down the results for the neural-text classifier over the different embedding types. Using only the hist embedding outperforms the phys-only setting for the symptoms cough, pain and fever. While the score for hist comes close to those for mean and concat, the latter usually still outperform the former for the normal notes, showing that there is some complementary information in the “history” and “physical examination” portions.
5 Discussion
Symptom extraction
Our analysis of simple tabular and textual baseline models revealed that the symptoms pain and fever are hardest to predict in both the tabular-only and the text-only setting. Combining both settings, i.e. integrating tabular background features in the extraction of the concepts from the text, and linking them through domain knowledge, may have potential for improved information extraction. Future work will focus on realizing this hybrid approach to improve upon the baseline results presented in Table 2.
Potential uses

### v3

@@ 2.2 Generating unstructured text with a large language model
Once the tabular patient record has been generated, we prompt a large language model (LLM, in our case GPT-4o) to write a clinical note based on the tabular variables associated with this fictional encounter. The scenario we simulate artificially is as follows. The patient goes to the primary care physician, telling them their symptoms and possible underlying conditions, along with additional context on the severity of these symptoms, when they started, and other details. The physician takes descriptive notes during this consultation, writing down the (recent) history prescribed by the patient. Then, based on the patient’s described complaints, they conduct a physical examination, writing down all findings. Both parts together then form the textual description of the patient encounter.
@@
To generate such a clinical note, we use a prompt which is generally structured like the example shown in Fig. 2. Additional example prompts can be found in Appendix C. Only the background variables and symptoms may be directly mentioned in the prompt, while the diagnoses pneumonia and common cold would not yet be known to a clinician who is taking notes during a consultation and are therefore left out. Realistically, they can still influence the content of the note through the descriptions of the symptoms that are included in the prompt, as will be explained later. The treatment and outcome are left out of the prompt as well, just like the non-clinical variables, as all of these are typically either unknown or irrelevant at the time of writing the note.We now describe the different parts of the prompt.
Presence of symptoms  The first block of information in the prompt concerns the symptoms experienced by the patient. We do not list the full set of symptoms exhaustively, since there is a possibility that a patient does not mention a symptom to the clinician, or that the clinician does not find it noteworthy enough to write down. We therefore ask our expert to list the probability of each symptom being mentioned in a clinical note, both when the symptom is positive and when it is negative. While we do not expect these arbitrary probabilities to generalize to all settings, it helps to bring some variety and realism in the notes we generate. The probabilities are as follows:
For each symptom, we sample whether it is to be mentioned in the prompt, conditional on its value, according to the probabilities stated above. As can be seen in Fig. 2, we explicitly tell the model what symptoms to mention and which to steer clear from. We randomly permute the ordering of the symptoms in each prompt.
Table 2: Descriptors used in the prompt to describe each symptom when it is present in the patient. Depending on the cause(s) of the symptom (as present in the tabular patient record), we randomly sample from a different set of descriptors.
Symptom Cause Descriptors
smoking during exercise, worse in morning, mild
hay fever light, mild, stuffy feeling, all closed up
pneumonia light, mild, severe, no clear cause
cough asthma attack-related, dry
smoking productive, mostly in morning, during exercise, gurgling
COPD phlegm, sputum, gurgling, worse when lying down
pain asthma tension behind sternum
COPD light, mild
Symptom descriptors To make the note realistic, the LLM must invent some context regarding the patient’s symptoms when writing the history portion of the note. We want this context to indirectly relate to the cause of these symptoms, as they would in a real patient encounter. For example, a cough induced by asthma would likely be momentarily and attack-related, while a cough resulting from pneumonia might be more persistent over the longer term. We therefore ask the expert to write down a list of adjectives or phrases describing each symptom, conditioned on the cause of the symptom. These descriptors can be found in Table 2. The list of possible causes for a symptom is simply the list of parents in the Bayesian network.
For each symptom which is present in the patient and selected to be mentioned in the note, we check the tabular patient record for the possible causes. For example, for the symptom cough, the possible causes are asthma, smoking, COPD, pneumonia and common cold (its parents in the Bayesian network). In the example in Fig. 2, asthma is the only cause present in the patient. We therefore randomly sample a descriptor from the list of descriptors for cough in the presence of asthma, in this case the adjective “dry”. This adjective is added in the prompt. If multiple causes are present, we find the strongest cause, and sample from that list. The strongest cause is pneumonia, followed by common cold, followed by all other causes. If neither pneumonia nor common cold is part of the multiple causes, we simply make a bag of all descriptors associated to the causes which are present in the patient, and sample from that bag. In the event that no causes are present, yet a symptom is still observed (which is possible due to the leak probabilities in the Noisy-OR distributions), we do not add a descriptor. Note that while the diagnoses pneumonia and common cold should not be mentioned explicitly, they indirectly and subtly influence the content of the note through the descriptors, adding another realistic dimension to the content of the note.
Underlying health conditions  While the diagnoses should not be mentioned directly in the note, it is realistic to assume that the note would mention underlying health conditions the patient may have. Since these health conditions are assumed to be known up-front, as they are part of the history of the patient, they may contribute a lot to the interpretation of the symptoms by both the patient themselves and the clinician writing down the note. We therefore add them to the prompt as well, as can be seen in Fig. 2. We do not force the LLM to explicitly mention these in the note, since it seems feasible that a clinician would not mention them every time. Should there be more than one underlying condition, we mention them all, randomly permuting the order. If there are no underlying health conditions, we simply remove this part of the prompt.
Additional instructions  We tell the LLM that the note must be structured with a “History” portion and a “Physical examination” portion. While the “History” portion describes the patient’s self-reported symptoms and underlying health conditions, which are in large part dictated by the prompt, the “Physical examination” portion leaves the LLM with more freedom to imagine additional clinical examinations which were performed on the patient. As such, the “Physical examination” portion has a lot of potential for adding complexity, clinical terminology and realism to the note.
We also add some additional instructions to the prompt, asking it not to mention any suspicions of possible diagnoses. We further tell the LLM it can imagine context or details, but no additional symptoms. We noticed that if we left this part out, the LLM would sometimes mention the symptoms we specifically asked to leave out. We ask not to mention patient gender or age, because our initial experiments without this instruction revealed that the LLM often used the same age and gender, which could confound or bias the notes. Finally, we add that the notes may be long (around 5 lines or more), to avoid the LLM being too succinct.
Special case: no respiratory symptoms  Around one third of all patients in our dataset (3629 out of 10,000 patients sampled from the BN) do not experience any respiratory symptoms at all (all symptoms are “no”). If we used the same prompt as before, this would result in an unrealistic clinical note, since the note would simply list all symptoms the patient does not have, without giving an actual reason for the patient’s visit. Furthermore, there would be little variation in these notes. An example is shown in Fig. 5 in Appendix B.
In these cases, it makes more sense to assume that the patient visits the doctor because of a complaint unrelated to the respiratory domain, such as back pain, stomach issues, a skin rash, etc. To generate these special cases, we use a special prompt, telling the LLM the patient does not experience any of the 5 respiratory symptoms. When the patient has at least one underlying health condition (which is the case in 239 out of 3629 special cases), we add this to the prompt in the same way as before, like the example in Fig. 6 in Appendix B. If not (i.e., for the remaining 3390 out of 3629 special cases), we tell the LLM not to mention any of those health conditions either, as shown in Fig. 7 in Appendix B. The latter prompt asks for three clinical notes at once, encouraging the LLM to be more creative and not repeat the same scenario every time, as well as being a little more cost-effective. This is possible because of the prompt being non-specific to any of these 3390 patients. We then randomly distribute all generated notes to each tabular patient record within this subset.
Consistency check  We add an additional filtering step to the generation process to better ensure the quality of the generated notes. We prompt the LLM once again, showing it the same information on the patient’s symptoms and underlying health conditions as in the previous prompt from Fig. 2. We then show the LLM the generated note, and ask it to decide whether it is consistent with this information or not. We ask it to provide an explanation if it deems the note inconsistent. This way, the LLM identified 52 out of $10{,}000$ notes that were inconsistent with the prompt. After manually verifying that this was indeed the case, we asked the LLM to regenerate these notes using the same prompt as before, afterwards finding that none of the newly generated notes were flagged as inconsistent anymore.
@@
Compact version  Real clinical notes can be challenging, often containing abbreviations, shortcuts and denser sentence structure. To make the notes more challenging, we ask the LLM to create a compact version of each note through an additional prompt, as can be seen in Fig. 2. Our dataset contains both the original note and the compact version of the note.
Prompting details  As a large language model, we opted for OpenAI’s GPT-4o model, using the version released in May 2024 [21]. We set the temperature to 1.2 to encourage some more variation in the notes. Before providing the case-specific prompt, we set the following system message: “You are a general practitioner, and need to summarize the patient encounter in a clinical note. Your notes are detailed and extensive.” We set the max_tokens parameter to 1000. All other parameters were set to their default values. Generating all 10,000 notes and their compact version cost around $150$\$.
2.3 Automated extraction of symptom spans
a LLM-based annotation of a normal clinical note with symptom spans. The model is prompted to return a JSON object of the form {‘‘symptom’’: symptom, ‘‘text’’: extracted phrase}. For clarity, we highlight the extracted spans in the note rather than displaying the raw JSON. Full task instructions are in Appendix D.
b LLM-based annotation of a compact clinical note using spans previously extracted from the corresponding normal note. Full task instructions are provided in Appendix D.
Figure 4: Prompting strategy for extracting symptom spans from (a) normal and (b) compact clinical notes using a large language model.
To further enhance the dataset, we annotated the notes with all spans that mention a symptom. We automatically extracted these phrases from the note using additional LLM calls.
Normal notes Fig. 4a shows the prompt and the extracted phrases for one of the normal notes. We first tell the LLM what symptoms the patient does and does not have, and then ask it to extract phrases that describe these symptoms. We use the system message to send extensive task instructions to the LLM, which we have included in Appendix D. We further process the extracted phrases by matching them to the note, disregarding capital letters and punctuation. If no exact match is found, we use another LLM call to correct the mistake, asking it to map the extracted phrase to an exact phrase in the text.
Compact notes For the compact notes, we proceed differently, as shown in Fig. 4b. We first tell the LLM which phrases were extracted from the normal note, and ask it to match these with the compact version of the note. We found that this worked better than asking it to extract phrases from the compact note from scratch. As with the normal notes, we post-process the extracted phrases by matching them to the text and asking the LLM to correct itself when no exact match is found.
Manual evaluation To assess the quality of the automated span extraction process, we conducted an evaluation study over 100 randomly selected notes. We manually annotated each of these notes with phrases that were missed and phrases that were extracted incorrectly. From this, we calculated precision (how many extracted phrases are correct, over the total number of extracted phrases per note) and recall (how many extracted phrases are correct, over the total number of phrases that should have been extracted per note). Averaging these metrics over this set of 100 notes yields a precision of $94.06\%$ and recall of $98.78\%$ for the normal notes, and a precision of $93.55\%$ and recall of $94.50\%$ for the compact notes.
Descriptive statistics The extracted spans allow us to calculate statistics on the presence of each of the 5 symptoms in the notes, which are reported in Table 3(b). We highlight some of the most interesting findings. Table 3(a) shows that on average, less spans were extracted for the compact notes than for the normal notes. The explanation is probably two-fold: while some descriptions could be missing from the compact notes entirely, their challenging nature also means more mistakes will be made in the automated extraction, leading to some spans being missed (as reflected by the lower average recall in our manual evaluation). The symptom pain is mentioned much less than the other 4 symptoms, even when it is present in the patient. Unsurprisingly, the average span length is much shorter in the compact notes, both in terms of words and characters. According to the extracted spans, most of the symptoms are described in the “history” portion of the clinical note, with proportions differing considerably between symptoms. The percentages of symptoms being mentioned at least once in the note (see Table 3(b)) correspond quite well with the probabilities used to build the prompt, as defined in Section 2.2. The average number of times a symptom is mentioned per note is much lower when the symptom is not present in the patient. In total, 16.3% of all normal notes do not mention any of the 5 symptoms at all (no spans extracted), while this percentage is 16.4% for the compact notes.
Table 3: Descriptive statistics on the symptoms mentioned in the notes, calculated via the extracted symptom spans.
(a) Statistics on the average number of spans per note, distribution of spans over symptoms, average length of spans, and location of spans (“history” vs. “physical examination” portion of the note). For each metric, we report normal / compact.
Dyspnea Cough Pain Nasal Fever All
Avg. # spans / note 0.96 / 0.90 1.03 / 0.96 0.52 / 0.49 0.85 / 0.81 0.79 / 0.74 4.16 / 3.90
Total % of spans 23.2% / 23.1% 24.9% / 24.6% 12.6% / 12.5% 20.4% / 20.7% 19.0% / 19.0% 100% / 100%
Avg. # words / span 8.34 / 4.86 8.83 / 5.22 9.34 / 5.96 8.92 / 5.98 7.04 / 4.07 8.46 / 5.17
Avg. # chars / span 57.21 / 30.45 56.70 / 32.14 63.54 / 36.57 63.05 / 41.74 44.82 / 23.95 56.71 / 32.73
% of spans in hist 79.3% / 78.8% 92.5% / 91.4% 81.6% / 80.8% 63.0% / 62.1% 65.9% / 65.2% 76.8% / 76.1%
% of spans in phys 20.7% / 21.2% 8.5% / 8.6% 18.4% / 19.2% 37.0% / 37.9% 34.1% / 34.8% 23.2% / 23.9%
(b) We show the percentage of notes where a symptom is mentioned (meaning there is at least one span mentioning the symptom in the note), and the average number of spans per note in which a symptom is mentioned. These are reported conditional on the value of the symptom (whether it is present in the patient or not).
Dyspnea Cough Pain Nasal Fever
yes no yes no yes no yes no high low none
% of notes where sympt. mentioned ($\geq$ 1 span / note) Normal Compact 99.1% 98.1% 66.5% 62.5% 97.0% 95.4% 64.5% 59.4% 81.7% 79.4 % 39.9% 37.1% 95.7% 95.4% 26.6% 25.1% 95.9% 95.7% 71.6% 71.4% 44.4% 42.4%
Avg. # spans where sympt. mentioned / note Normal Compact 2.03 1.91 0.70 0.65 1.78 1.66 0.65 0.60 1.28 1.21 0.40 0.37 2.47 2.38 0.30 0.28 2.46 2.33 1.55 1.48 0.52 0.48
3 Utility and discussion
We now present a series of experiments to illustrate the utility and potential of our dataset. First, Section 3.1 reports the results of an expert evaluation study assessing the quality of the artificial clinical notes generated by the LLM. Next, Section 3.2 evaluates the performance of simple symptom prediction models applied to both the tabular and textual portions of the dataset. Finally, Section 3.3 explores the broader research applications of the dataset, including but not limited to clinical information extraction.
3.1 Expert evaluation
@@
    Realism: The context and details invented by the LLM and added to the “history” portion of the note should be realistic given the symptoms experienced by the patient, as detailed in the prompt. Furthermore, the elements that are checked in the “physical examination” must be realistic in light of the patient information described in the “history”. Both aspects are evaluated separately, focusing on content (what information is included in the note), rather than format (how the information is written down).
@@
Normal Compact
mean 4.69 4.53 4.15 4.92 4.88 4.02
std 0.12 0.21 0.30 0.07 0.10 0.31
Results  Our artificial notes were rated as highly consistent with the prompt, and therefore with the information present in the tabular portion of the dataset. As shown in Appendix E.2, a large majority of the inconsistencies arise due to a violation of the additional instructions (usually by inventing additional symptoms), while the key information included in the prompt was still conveyed correctly in the note. High consistency between the tabular variables and the concepts mentioned in the text makes our dataset a reliable resource for information extraction tasks.
@@
3.2 Symptom predictor baselines
In order to set a baseline for future information extraction tasks, we run various prediction models on both the tabular and textual parts of the dataset. These models are trained to predict each of the five symptoms: dyspnea, cough, pain, fever and nasal. Two of our baselines only get to see the tabular portion of the dataset at the input: Bayesian network (BN-tab) and XGBoost (XGBoost-tab). We use these models to predict each symptom in three settings, differing from one another in the set of tabular features that are taken as an input, which we call the evidence:
@@ 3.2 Symptom predictor baselines
Apart from the tabular-only baselines, we also train baselines that get to see the text. Our neural-text classifier takes only the text as an input (in the form of a pretrained clinical sentence embedding) and outputs the probability that a symptom is mentioned in the text. We extend this text-only baseline by concatenating a numerical representation of the tabular features to the text embedding at the input, forming the neural-text-tab baseline. Again, we do this for each of the three evidence settings outlined above. Note that this is the only model that combines both the background knowledge available in the tabular features with the unstructured text, and it does so in a naive way. Future work will focus on improving the performance of this model by exploiting the relations between the tabular concepts, as envisioned in Fig. 1.
Section 3.2.1 provides a detailed overview of each baseline model, followed by Section 3.2.2, which analyzes their performance on the task of automated symptom extraction using the SimSUM dataset.
3.2.1 Models
BN-tab We provide the causal structure in Fig. 2 to the BN, and learn all parameters in the conditional probability tables (CPTs), Noisy-OR distributions, logistic regression model and Poisson regression model from the training data. In each case, we use maximum likelihood estimation to estimate the parameters. Where we don’t directly learn a CPT (for the variables dyspnea, cough, pain, nasal, antibiotics and #days), we evaluate the learned distribution for each combination of child and parent values to obtain a CPT. For more details, we refer to Appendix F.1. We then use variable elimination over the full joint distribution to evaluate the capability of the learned BN to predict each of the symptoms, taking different sets of variables as evidence according to the three settings described earlier (all, no-sympt and realistic).
XGBoost-tab We train an XGBoost classifier for each symptom in combination with each of the three evidence settings, meaning each classifier sees a different set of tabular features at the input. We optimize the hyperparameters separately for each combination (15 in total) using 5-fold cross-validation. For more details, we refer to Appendix F.2.
Neural-text We train a neural classifier that takes only the text as an input and is trained to predict the probability a symptom is mentioned. We train separate classifiers for each symptom. We first split the text into sentences, and transform these into an embedding using the pretrained clinical representation model BioLORD-2023 [26]. We explore four settings for turning these sentence embeddings into a single note embedding:
Additionally, we create symptom-specific span embeddings as well. To this end, we take the set of spans mentioning the symptom (as extracted automatically in Section 2.3), transform each of these into their BioLORD embedding, and take the average to get one span embedding per symptom.
No matter the type, each note embedding is fed into a multi-layer perceptron with one hidden layer, followed by a Sigmoid activation for the symptoms dyspnea, cough, pain and nasal, and a Softmax activation with 3 outputs heads for the symptom $fever$. We trained each model (i.e. each combination of symptom and embedding type) using the binary or multiclass cross-entropy objective over the symptom labels. For more details, we refer to Appendix F.3.
Neural-text-tab  We extend the neural-text baseline by concatenating the mean text embeddings with the tabular variables at the input of each neural classifier. All categorical tabular variables were first transformed to a one-hot encoding, while the variable #days was preprocessed using standard scaling. We used the same architecture as the neural-text baseline (only changing the dimension of the input layer), and again trained separate classifiers for each symptom combined with each evidence setting (all, no-sympt and realistic). For more details, we refer to Appendix F.4.
Setup  We use a random $8{,}000$/$2{,}000$ split to obtain a train and test subset of SimSUM. We use cross-validation on the train set to tune any hyperparameters associated with each model, and report the average F1-score over the test set after training with 5 different seeds [^2] Seeding randomizes the initialization of the model weights, as well as batching.. For the binary symptoms, we use a $0.5$ decision threshold. For fever, which has three possible categories, the class with the highest predicted probability is chosen.
3.2.2 Results
Table 5: F1-score obtained over the test set for each of our baseline models. For fever, we report the macro F1 score over all three categories. The results for the text classifiers trained over the normal vs. the compact version of the notes are grouped together for readability. We report results for the mean embedding type, while neural-text baseline results for the other embedding types can be found in Table 6. We underline the best result obtained by the tabular-only models, while the best overall result per symptom and note version is in shown in bold.
@@ 3.2.2 Results
- normal + all 0.9544 0.9555 0.8191 0.9613 0.9091
- normal + no-sympt 0.9594 0.9577 0.8155 0.9611 0.8996
- normal + realistic 0.9569 0.9596 0.8266 0.9591 0.9029
@@
- compact + all 0.9532 0.9395 0.7873 0.9589 0.8962
- compact + no-sympt 0.9429 0.9456 0.7856 0.9632 0.8973
- compact + realistic 0.9379 0.9480 0.7922 0.9630 0.9025
Table 6: Results for the neural-text baseline using different embedding types at the input. We report F1-score over the test set for both the normal and compact versions of the notes. The best results per symptoms and per note version is shown in bold, the runner-up is underlined. Fig. 10 in the Appendix compares these results visually.
Normal dyspnea cough pain nasal fever
Sentence embedding
- hist 0.9520 0.9652 0.8089 0.9504 0.9040
- phys 0.8969 0.8226 0.6956 0.9565 0.8523
- mean 0.9617 0.9603 0.8143 0.9628 0.9096
- concat 0.9504 0.9572 0.8268 0.9589 0.9044
Span embedding 0.9730 0.9785 0.8731 0.9712 0.9226
Compact dyspnea cough pain nasal fever
Sentence embedding
- hist 0.9307 0.9523 0.7734 0.9515 0.9001
- phys 0.8824 0.7802 0.6514 0.9531 0.8315
- mean 0.9444 0.9397 0.7940 0.9622 0.9010
- concat 0.9522 0.9473 0.7934 0.9411 0.9007
Span embedding 0.9685 0.9719 0.8613 0.9705 0.9207
@@
Table 5 compares the results obtained for all baselines. The tabular-only baselines (BN-tab and XGBoost-tab) perform consistently worse than the baselines that include text (neural-text and neural-text-tab). The evidence setting where all other features are included as evidence usually performs better than the other settings.
@@
For the text baselines, we note that there is little room for improvement in the dyspnea, cough and nasal classifiers. On the other hand, the symptoms pain and fever are harder to predict. We can assume that this is largely because these symptoms are mentioned less often in the notes, even if they are present in the patient, as supported by Table 3(b). We also see a consistent gap in performance between the normal and compact notes, which can be attributed to the higher complexity of the latter. While the neural-text-tab baseline does not manage to outperform the neural-text baseline for the normal notes, we do see marginal improvements by including the tabular features in the case of the compact notes.
Table 6 further breaks down the results for the neural-text classifier over the different embedding types. We clearly see that the span embeddings perform best, since this method preselects phrases that are important for the prediction of a certain symptom. While the same phrases are still implicitly embedded in the other types, the important information gets diluted as sentence embeddings are averaged, and might even be lost entirely. However, in a realistic setting, one would rarely have access to these symptom-specific span embeddings, as this would require either a manual annotation effort, or a (small) local LLM to facilitate automated annotation in a way that preserves patient privacy. We do not expect such a local LLM to achieve near the same span detection performance as GPT-4o, which we were only able to use in this case due to the artificial nature of our notes. The performance achieved with these span embeddings should therefore be seen as an upper bound for our benchmark.
@@
Comparing the four different sentence embedding types, the performance difference between the mean and concat settings is usually small, with mean slightly outperforming concat overall. While the score for hist comes close to those for mean and concat, the latter two usually still outperform the former for the normal notes, showing that there is some complementary information in the “history” and “physical examination” portions. At the same time, we note that using only the hist embedding clearly outperforms the phys-only setting. This makes sense, as the “history” section of the note outlines the symptoms experienced by the patient more clearly.
3.3 Intended uses
The SimSUM dataset is primarily designed to support research on clinical information extraction in settings where tabular background variables are available. A key objective for future work is to realize the approach illustrated in Fig. 1, in which domain knowledge links structured features to textual concepts, enabling more accurate information extraction. As demonstrated in the previous section, even a simple method—concatenating tabular features with text embeddings at the input of a neural classifier—already improves symptom extraction performance, particularly on the more challenging, compact notes. This suggests that background information can help recover symptoms that are mentioned less explicitly in the text, underscoring the utility of our dataset for developing and evaluating such approaches.